OpenStudy (anonymous):

find the difference (12s^2+8s-6)-(9s^2-2s+5).

OpenStudy (anonymous):

We have : \[12s^2+8s-6-(9s^2-2s+5)=12s^2+8s-6-9s^2+2s-5\\ =(12-9)s^2+(8+2)s-6-5\\ =3s^2+10s-11 \]

OpenStudy (anonymous):

thank you so much